# Wall crossing and the Fourier–Mukai transform for Grassmann flops

### v1

@@ 3.2.3. The non-abelian quotients
$N_{\beta,\vec{d}}=\prod_{j=1}^{r}$ $\left(\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}\right)$
(3.2.5) $\cdot$ $\prod_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{r}\frac{\prod_{h=-\infty}^{d_{i}-d_{j}}(y_{i}-y_{j}+hz)}{\prod_{h=-\infty}^{0}(y_{i}-y_{j}+hz)}.$
@@
$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\ldots,-e_{r}}$
$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\ldots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\ldots,d_{r}}.$
@@
$I_{X_{\pm}}(Q,q_{\pm},z)=I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,\vec{q}_{+},z)|_{q_{1,\pm}=\cdots=q_{s,\pm}=q_{\pm}}.$
The following proposition is a corollary of the work of Coates–Lutz–Shafi in [CLS22] on the abelian/non-abelian correspondence. One could also prove it directly by generalizing the work of Oh in [Oh21] on $I$-functions of Grassmann bundles.
@@
As described in [CLS22], $I_{X_{\pm}}$ is the Givental–Martin modification of $I_{X_{T,\pm}}$, so the proposition essentially follows from [CLS22, Theorem 1.11]. More precisely, we apply the argument of Theorem 5.11 of [CLS22], which provides an abelian/non-abelian correspondence for twisted theories over a flag bundle. Because that theorem is not stated in exactly the generality which we need here, we summarize the argument below and explain the necessary adjustments to the case at hand. We will prove the proposition for $X_{+}$.
@@
$Z:=\operatorname{Gr}(r,E)=\operatorname{tot}\left(\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}G.$
@@
$Z_{T}:=\operatorname{tot}\left(\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}T=\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E).$
Let $I_{Z_{T}}$ denote the Brown $I$-function for $Z_{T}$ as in [Bro14], and let $I_{Z_{T}}^{e_{\mathbb{T}}^{-1}((S_{T}\otimes p^{*}F^{\vee}))}$ denote the corresponding $((S_{T}\otimes p^{*}F^{\vee}),e^{-1}_{\mathbb{T}}(-))$-twisted $I$-function, where we use $S_{T}$ to denote the vector bundle on $Z_{T}$ induced by the right standard representation $R$ from (2.1.1) restricted to $T\subset G$. Then $I_{Z_{T}}^{e_{\mathbb{T}}^{-1}((S_{T}\otimes p^{*}F^{\vee}))}$ is exactly the $I$-function $I_{X_{T,+}}$ from the previous section.
Following the notation and discussion in Section 5.2 of [CLS22] (in particular Theorem 5.11), we apply a further twist to $I_{X_{T,+}}=I_{Z_{T}}^{e^{-1}_{\mathbb{T}}((S_{T}\otimes p^{*}F^{\vee})}$ by $(\Phi,\mathbf{c}^{\prime})$ where $\Phi\to Z_{T}$ is the root bundle with respect to $G$ and $\mathbf{c}^{\prime}$ is the $\mathbb{C}^{*}$-equivariant Euler class with respect to the $\mathbb{C}^{*}$-action scaling the fibers of $\Phi.$ Denote the equivariant parameter of this $\mathbb{C}^{*}$ action by $\lambda$. Specializing variables to $q_{1,+}=\cdots=q_{s,+}=q_{+},$ projecting via the map on Givental spaces defined in [CLS22]
$p:\mathcal{H}_{Z_{T}}^{W}\to\mathcal{H}_{Z},$
@@
By the proof of [CLS22, Theorem 5.11], this function lies in the $((S_{+}\otimes p^{*}F^{\vee}),e^{-1}_{\mathbb{T}})$-twisted Lagrangian cone of $Z=\operatorname{Gr}(r,E)$. While the theorem in loc. cit. is only stated for the twisted theory $(L,e_{\mathbb{C}^{*}}(-))$, where $L$ is a line bundle and $e_{\mathbb{C}^{*}}(-)$ is an equivariant Euler class, by [CLS22, Remark 3.7] the result holds for other multiplicative classes as well. The further extension to the case that $L$ is replaced by the sum of line bundles also follows the same argument.
By definition, the $((S_{+}\otimes p^{*}F^{\vee}),e^{-1}_{\mathbb{T}})$-twisted Lagrangian cone of $\operatorname{Gr}(r,E)$ is equal to the $\mathbb{T}$-equivariant Lagrangian cone of $X_{+}$.
4. Analytic continuation: the abelian case
In this section we explicitly compute the analytic continuation of the $I$-functions of the associated abelian quotients $X_{T,\pm}$ and show that they are related by symplectic transformation. This was proven in [CIJ18] in the case $B$ is a point.
4.1. Projective bundles
@@
We begin by writing $I_{X_{\pm}}$ as $\psi(H_{X_{\pm}})$ where $\psi$ is as in (3.1.1). The relative Euler sequence for $X_{\pm}$ is
$0\to\mathcal{O}_{X_{\pm}}\to R\otimes p^{*}E\oplus R^{\vee}\otimes p^{*}F^{\vee}\to TX_{\pm}/B\to 0$
where $R$ is the vector bundle defined in Section 2.1. Because the Gamma class is multiplicative, we have $\hat{\Gamma}$, we have that
$\hat{\Gamma}_{X_{\pm}}=$ $\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)$
$=$ $\hat{\Gamma}_{B}\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y))\prod_{i=1}^{n}\Gamma(1+(z_{i}+y)).$
Note that the hypergeometric modification (3.2.3) is equal to
(4.1.1) $\frac{1}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}\prod_{j=1}^{n}\frac{\Gamma(1+(-x_{j}-y)/z)\Gamma(1+(z_{j}+y)/z)}{\Gamma(1+(-x_{j}-y)/z+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/z-\beta\cdot M_{j}+d)}$
@@
$H^{B}$ $:=(2\pi\mathbbm{i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}$
$=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}(TB\oplus E\oplus F^{\vee})$ and
$\mu_{X}(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\right)\phi.$
Define the $H^{*}_{\mathbb{T}}(X_{+})$-valued function
(4.1.2) $H_{X_{+},\beta}(q_{+})=$
$\sum_{e\in\mathbb{Z}}\frac{q_{+}^{-y/2\pi\mathbbm{i}+e}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi\mathbbm{i}+\beta\cdot L_{j}+e)\Gamma(1+(z_{j}+y)/2\pi\mathbbm{i}-\beta\cdot M_{j}-e)}.$
@@
(4.1.3) $H_{X_{-},\beta}(q_{-})=$
$\sum_{d\in\mathbb{Z}}\frac{q_{-}^{y/2\pi\mathbbm{i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/2\pi\mathbbm{i}-\beta\cdot M_{j}+d)}.$
@@
(4.1.4) $I_{X_{\pm}}(Q,q_{\pm},z)$ $=\psi\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{\pm},\beta}\right).$
We will fix $\beta$ and relate $H_{X_{+},\beta}$ and $H_{X_{-},\beta}$ via analytic continuation and symplectic transformation. Recall Definition 2.8 in the special case $r=1$. In this case $\mathbb{U}_{H}:H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{B_{k}|X_{-}})\in H^{*}_{\mathbb{T}}(B_{k}^{-})$ to
(4.1.5) $\sum_{l=1}^{n}e^{(n-1)\cdot(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2\mathbbm{i})}\phi_{kl}(\alpha)/e_{\mathbb{T}}(N_{B_{l}|X_{+}})$
where $\phi_{kl}:H^{*}_{\mathbb{T}}(B_{k}^{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(B_{l}^{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{k}^{+}:B_{k}^{+}\to B$ and $\pi_{l}^{-}:B_{l}^{-}\to B$. By restricting $H_{X_{+},\beta}$ to the fixed points of the $\mathbb{T}$-action and using Mellin–Barnes integrals, we may analytically continue $H_{X_{+},\beta}$ to $q_{+}=\infty.$ Because the method of computation is by now somewhat standard (see e.g. [CIJ18]), we relegate the proof to Appendix A. We obtain the following.
@@
(4.1.6) $\mathbb{U}_{H}H_{X_{-},\beta}=\widetilde{H_{X_{+},\beta}}$
@@
${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\mathbb{U}}$
@@
The proof of the equality is immediate from Definition 3.5, (4.1.4) and (4.1.6). The fact that $\mathbb{U}$ is symplectic and has a well-defined non-equivariant limit is Corollary 3.6. The commuting diagram follows from Proposition 2.9. ∎
4.2. Products of projective bundles

### v2

@@ 3.2.3 The non-abelian quotients
$N_{\beta,\vec{d}}=\prod_{j=1}^{r}\Biggl{(}\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}\Biggr{)}$
×${}\times\prod_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{r}\frac{\prod_{h=-\infty}^{d_{i}-d_{j}}(y_{i}-y_{j}+hz)}{\prod_{h=-\infty}^{0}(y_{i}-y_{j}+hz)}.$ (3.4)
@@
$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\dots,-e_{r}},$
$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\dots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\dots,d_{r}}.$
@@
$I_{X_{\pm}}(Q,q_{\pm},z)=I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,\vec{q}_{\pm},z)|_{q_{1,\pm}=\cdots=q_{s,\pm}=q_{\pm}}.$
The following proposition is a corollary of Coates–Lutz–Shafi in [17] on the abelian/non-abelian correspondence. One could also prove it directly by generalizing the work of Oh in [29] on $I$-functions of Grassmann bundles.
@@
As described in [17], $I_{X_{\pm}}$ is the Givental–Martin modification of $I_{X_{T,\pm}}$, so the proposition essentially follows from [17, Theorem 1.11]. More precisely, we apply the argument of [17, Theorem 5.11], which provides an abelian/non-abelian correspondence for twisted theories over a flag bundle. Because that theorem is not stated in exactly the generality which we need here, we summarize the argument below and explain the necessary adjustments to the case at hand. We will prove the proposition for $X_{+}$.
@@
$Z:=\operatorname{Gr}(r,E)=\operatorname{tot}\left(\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}G.$
@@
$Z_{T}:=\operatorname{tot}\left(\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}T=\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E).$
Let $I_{Z_{T}}$ denote the Brown $I$-function for $Z_{T}$ as in [7], and let $I_{Z_{T}}^{e_{\mathbb{T}}^{-1}((S_{T}\otimes p^{*}F^{\vee}))}$ denote the corresponding $\bigl{(}\bigl{(}S_{T}\otimes p^{*}F^{\vee}\bigr{)},{\rm e}^{-1}_{\mathbb{T}}(-)\bigr{)}$-twisted $I$-function, where we use $S_{T}$ to denote the vector bundle on $Z_{T}$ induced by the right standard representation $R$ from (2.1) restricted to $T\subset G$. Then $I_{Z_{T}}^{e_{\mathbb{T}}^{-1}((S_{T}\otimes p^{*}F^{\vee}))}$ is exactly the $I$-function $I_{X_{T,+}}$ from the previous section.
Following the notation and discussion in [17, Section 5.2] (in particular, Theorem 5.11), we apply a further twist to
$I_{X_{T,+}}=I_{Z_{T}}^{{\rm e}^{-1}_{\mathbb{T}}((S_{T}\otimes p^{*}F^{\vee})}$
by $(\Phi,\mathbf{c}^{\prime})$ where $\Phi\to Z_{T}$ is the root bundle with respect to $G$ and $\mathbf{c}^{\prime}$ is the $\mathbb{C}^{*}$-equivariant Euler class with respect to the $\mathbb{C}^{*}$-action scaling the fibers of $\Phi$. Denote the equivariant parameter of this $\mathbb{C}^{*}$ action by $\lambda$. Specializing variables to $q_{1,+}=\cdots=q_{s,+}=q_{+},$ projecting via the map on Givental spaces defined in [17]
$p\colon\ \mathcal{H}_{Z_{T}}^{W}\to\mathcal{H}_{Z},$
@@
By the proof of [17, Theorem 5.11], this function lies in the $\bigl{(}\bigl{(}S_{+}\otimes p^{*}F^{\vee}\bigr{)},{\rm e}^{-1}_{\mathbb{T}}\bigr{)}$-twisted Lagrangian cone of $Z=\operatorname{Gr}(r,E)$. While the theorem in loc. cit. is only stated for the twisted theory $(L,e_{\mathbb{C}^{*}}(-))$, where $L$ is a line bundle and $e_{\mathbb{C}^{*}}(-)$ is an equivariant Euler class, by [17, Remark 3.7] the result holds for other multiplicative classes as well. The further extension to the case that $L$ is replaced by the sum of line bundles also follows the same argument.
By definition, the $\bigl{(}\bigl{(}S_{+}\otimes p^{*}F^{\vee}\bigr{)},{\rm e}^{-1}_{\mathbb{T}}\bigr{)}$-twisted Lagrangian cone of $\operatorname{Gr}(r,E)$ is equal to the $\mathbb{T}$-equivariant Lagrangian cone of $X_{+}$. ∎
4 Analytic continuation: The abelian case
In this section, we explicitly compute the analytic continuation of the $I$-functions of the associated abelian quotients $X_{T,\pm}$ and show that they are related by symplectic transformation. This was proven in [14] in the case $B$ is a point.
4.1 Projective bundles
@@
We begin by writing $I_{X_{\pm}}$ as $\psi(H_{X_{\pm}})$ where $\psi$ is as in (3.1). By (2.5) in the case $r=1$ we obtain the relative Euler sequence
$0\to\mathcal{O}_{X_{-}}\to S_{-}\otimes p^{*}E\oplus S_{-}^{\vee}\otimes p^{*}F^{\vee}\to TX_{-}/B\to 0$
and similarly for $X_{+}$. Because the Gamma class is multiplicative, we have that
$\hat{\Gamma}_{X_{\pm}}=\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)=\hat{\Gamma}_{B}\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y))\prod_{i=1}^{n}\Gamma(1+(z_{i}+y)).$
Note that the hypergeometric modification (3.3) is equal to
$\frac{1}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}\prod_{j=1}^{n}\frac{\Gamma(1+(-x_{j}-y)/z)\Gamma(1+(z_{j}+y)/z)}{\Gamma(1+(-x_{j}-y)/z+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/z-\beta\cdot M_{j}+d)}.$
@@
$H^{B}:=(2\pi{\rm i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}\bigl{(}TB\oplus E\oplus F^{\vee}\bigr{)}$ and
$\mu_{X}(\phi)=\biggl{(}\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\biggr{)}\phi.$
Define the $H^{*}_{\mathbb{T}}\bigl{(}X_{+}\bigr{)}$-valued function
$H_{X_{+},\beta}(q_{+})$
$\qquad{}=\sum_{e\in\mathbb{Z}}\frac{q_{+}^{-y/2\pi{\rm i}+e}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi{\rm i}+\beta\cdot L_{j}+e)\Gamma(1+(z_{j}+y)/2\pi{\rm i}-\beta\cdot M_{j}-e)}.$
@@
$H_{X_{-},\beta}(q_{-})$ (4.1)
$\qquad{}=\sum_{d\in\mathbb{Z}}\frac{q_{-}^{y/2\pi{\rm i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi{\rm i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/2\pi{\rm i}-\beta\cdot M_{j}+d)}.$
@@
$I_{X_{\pm}}(Q,q_{\pm},z)$ ${}=\psi\biggl{(}\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{\pm},\beta}\biggr{)}.$ (4.2)
We will fix $\beta$ and relate $H_{X_{+},\beta}$ and $H_{X_{-},\beta}$ via analytic continuation and symplectic transformation. Recall Definition 2.9 in the special case $r=1$. In this case, $\mathbb{U}_{H}\colon H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{B_{k}|X_{-}})\in H^{*}_{\mathbb{T}}(B_{k}^{-})$ to
$\sum_{l=1}^{n}{\rm e}^{(n-1)\cdot(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2{\rm i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2{\rm i})}\phi_{kl}(\alpha)/e_{\mathbb{T}}(N_{B_{l}|X_{+}}),$ (4.3)
where $\phi_{kl}\colon H^{*}_{\mathbb{T}}(B_{k}^{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(B_{l}^{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{k}^{+}\colon B_{k}^{+}\to B$ and $\pi_{l}^{-}\colon B_{l}^{-}\to B$. By restricting $H_{X_{+},\beta}$ to the fixed points of the $\mathbb{T}$-action and using Mellin–Barnes integrals, we may analytically continue $H_{X_{+},\beta}$ to $q_{+}=\infty.$ Because the method of computation is by now somewhat standard (see, e.g., [14]), we relegate the proof to Appendix A. We obtain the following.
@@
$\mathbb{U}_{H}H_{X_{-},\beta}=\widetilde{H_{X_{+},\beta}}$ (4.4)
@@
${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\mathbb{U}}$
@@
The proof of the equality is immediate from Definition 3.5, (4.2) and (4.4). The fact that $\mathbb{U}$ is symplectic and has a well-defined non-equivariant limit is Corollary 3.6. The commuting diagram follows from Proposition 2.10. ∎
4.2 Products of projective bundles